OpenStudy (anonymous):

if f(x)=e^( sin x), how many zeros does f'(x) have on the interval [ 0, 2pi] ? I used chain rule to find the first derivative and set it equal to zero, but I'm not sure where to go from there. Help ..please!

OpenStudy (e.mccormick):

Well, if you got f' and set it to 0, you have a periodic answer. You need to look at that solution in the domain that was supplied of 0 to 2pi.

OpenStudy (anonymous):

Dude your く isn't showing!

OpenStudy (e.mccormick):

From what you said, you have \[f(x)=e^{\sin x} \rightarrow f'(x)=(\cos x )e^{\sin x}\] Well, that means it is 0 when either factor is 0. So on [0, 2pi] when is either cos x or e^(sin x) 0?

OpenStudy (e.mccormick):

@wio Yah, I need to re-edit so the number does not kill it. Nihon go ga wakarimasuka

OpenStudy (anonymous):

少し分かりますが、単語が苦手です。

OpenStudy (anonymous):

Basically, since \(e^{g(x)}=0\) is never the case, you only need to look at \(g'(x)\) in this case.

OpenStudy (e.mccormick):

Yah, which makes it simpler because he only has cos x to deal with. 漢字は難しいでしょ

OpenStudy (anonymous):

Oh! ok, got it. so then, there's two zeros in the interval.

OpenStudy (e.mccormick):

If that is inclusive, rather than [0, 2pi), then...

OpenStudy (anonymous):

ひらがなをつかってほしかったら、そうしますが、漢字ができるのでぼくはすきです。

OpenStudy (e.mccormick):

Actually, yah. I was thinking sin x for a sec. cos x will only be when you are at the top and the bottom. But it does point out where you need to be careful.

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

@ Wio ..I don't actually understand what you are saying.

OpenStudy (e.mccormick):

He and I are speaking in Japanese, completely unrelated to your problem. LOL

OpenStudy (anonymous):

I figured it was something like that.

OpenStudy (anonymous):

lol

OpenStudy (e.mccormick):

He liked my icon, which is Jaopanese for Mathematics, and we have traded a bit on how it can be a challenege to learn, or basically that.

OpenStudy (e.mccormick):

And half of that is subtext rather than text because you would only comment on the ku being missing if you knew it and liked it.